What are the zeros of f(x)=x-40+6x^2 ? I think the answer is (-16,15) but I'm not sure.
Answer by scott8148(6628)

6x^2+x-40 factors into (3x+8)(2x-5)

if (3x+8)(2x-5)=0, then x=-8/3 and x=5/2